A data set lists the weights​ (grams) of a type of coin. Those weights have a mean of 5.41478 g and a standard deviation of 0.06439 g. Identify the weights that are significantly low or significantly high.


1
Expert's answer
2021-09-21T02:44:38-0400

QUESTION

Consider a value to be significantly low if its z score is less than or equal to -2 or consider the value to be significantly high if its z score is greater than or equal to 2. A data set lists the weights​ (grams) of a type of coin. Those weights have a mean of 5.41478 g and a standard deviation of 0.06439 g. Identify the weights that are significantly low or significantly high.

SOLUION

Given, mean, "\mu=5.41478" and standard deviation, "\sigma=0.06439"

The formula to z-score, "Z=\frac{x-\mu}{\sigma}"

But the values of "Z" are given. So, "x=\mu+Z\sigma"

We are going to calculate "x" when "Z" is at "-2" and "2"

When "Z" is at "-2",

Then "x=5.41478-2(0.06439)"

"x=5.286"

When "Z" is at "2"

Then "x=5.41478+2(0.06439)"

"x=5.54356"

ANSWER: The weight that are significantly low "=5.286" and the weight that is significantly high "=5.54356"